find the area of a circle with a circumference of 20(pie)
*pie= 3.14

Answer by nerdybill(5404)

find the area of a circle with a circumference of 20(pi)
.
Since the equation for circumference is:
Circumference = 2(pi)r
.
And, since they told us circumference is 20(pi) we now have:
20(pi) = 2(pi)r
20 = 2r
10 = r
.
Area of a circle is (pi)r^2
Therefore:
area = (pi)(10)^2
area = (3.14)(10)^2
area = (3.14)(100)
area = 314 square units